OpenStudy (anonymous):

Solve the equation for the letter y: -2x + 4y = 8

OpenStudy (anonymous):

y= 1/2x+2

OpenStudy (hoblos):

-2x + 4y = 8 add 2x 4y = 2x+8 divide by 4 y= (1/2)x +2